OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

The equation is \[f(x) = 3(x - 11)^{2} + 2\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

The average rate of change is the ratio of the change in the function value to the change in the independent variable over the interval, i.e. \(\dfrac{f(b)-f(a)}{b-a}\). Here \(a=4\) and \(b=7\).